# Comparison of chiral limit studies in curvature mass versus on-shell renormalized quark-meson model using ChPT

### v1

@@ III.1 QMVT versus RQM model order parameters and its derivatives in the path to chiral limit
It is worthwhile to compare the $\beta_{cep}$ for the case of the $m_{\sigma}=400$ MeV, though the order parameter and its derivative plots are not shown. The critical end point fraction in the QMVT model is found at the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{QMVT})}=0.2215\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(30.57,~{}109.88)$ MeV when the $m_{\sigma}=400$ MeV while the similar cep in the RQM model is found at the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{RQM})}=0.3795\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(52.38,~{}188.25)$ MeV. It is known that the strength of the chiral transition becomes stronger for smaller scalar $\sigma$ meson masses. The above feature gets verified as we are finding the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{QMVT})}=0.2215>\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{QMVT})}=0.2028$ in the QMVT model while we find the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{RQM})}=0.3795>\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{RQM})}=0.3642$ for the RQM model.
[…]
@@ III.2 Comparing the $m_{\sigma}$ towards the chiral limit
Note that the curvature and pole masses are same in the QMVT model whereas the curvature mass of the $\pi(\sigma)$ meson $m_{\pi,c}(m_{\sigma,c})=135.98(376.74)\text{ MeV}$ is smaller (significantly smaller) than its pole mass of $m_{\pi}(m_{\sigma})=138(530)\text{ MeV}$ for $\beta=1$ in the RQM model due to the on-shell renormalization of the parameters. The QMVT model temperature variation of the $\sigma$ mass in the Fig. 3, begins from the same value $m_{\sigma}=530$ MeV for all the $\beta$ fractions whereas the RQM model curvature mass of the $\sigma$ meson in the Fig. 3 starts from the $m_{\sigma,c}=376.74,420.29,443.01\text{ and }474.75$ MeV respectively for the $\beta=1.0,0.5,0.3642\text{ and }0.15$. Thus we find that the vacuum curvature mass of the $\sigma$ meson in the RQM model, is lowest at the physical point and it increases successively when one moves towards the chiral limit after reducing the ($m_{\pi},m_{K}$) by a fraction $\beta$ of their physical point value. The behavior of the $m_{\sigma}$ in our chiral limit study for both the QMVT and RQM models, stands in contrast to its behavior in the e-MFA-FRG, QM model study of the Ref.Resch where in the aftermath of the successive increase in the scale Λ=700,1000,1100 and $\Lambda=700,1000,1100\text{ and }1143$ corresponding to the successive decrease of the source strengths $j_{x}\text{ and }j_{y}$ governed by the $\alpha=\sqrt{\beta}=1.0,0.17,0.04\text{ and }0$, the $\sigma$ mass registers a significant successive decrease towards the chiral limit as evident from the Fig.2(c) of the Ref.Resch . Here, it is relevant to point out that the $\sigma$ mass in the e-MFA-FRG study is calculated from the curvature of the effective potential that gets evaluated by the FRG flow equation. Recall that the curvature masses of mesons are used for fixing the parameters in the QMVT model and this context of incorporating quark one-loop vacuum correction is analogous to that of the e-MFA-FRG study in the Ref.Resch where the vacuum value of the $m_{\sigma}$ decreases significantly towards the chiral limit while the QMVT model $m_{\sigma}$ does not change towards the chiral limit. In complete contrast of the above two scenarios, being lowest at the physical point and different from its pole mass $m_{\sigma}=530$ MeV, the vacuum curvature mass $m_{\sigma,c}$ in the on-shell renormalized RQM model increases towards the chiral limit.
@@
Figure 4:  Columbia plot with the $U_{A}$(1) anomaly is presented in the left (right) panel (a) ((b)) for the QMVT (RQM) model when $m_{\sigma}=400$ MeV. The $\mu-m_{K}$ ($m_{\pi}-m_{K}$) plane presents the chiral transition for the $m_{\pi}=0$ ($\mu=0$). The second order $Z(2)$ critical solid black line demarcating the crossover from the first order region, intersects the SU(3) symmetric $m_{\pi}=m_{K}$ green dash line respectively at the critical pion mass of $m_{\pi}^{c}=78.45$ and $m_{\pi}^{c}=134.16$ MeV in (a) for the QMVT model and (b) for the RQM model. It terminates at the terminal pion mass of $m_{\pi}^{t}=98.15$ MeV in (a) for the QMVT model and $m_{\pi}^{t}=168.39$ MeV in (b) for the RQM model. Solid blue line of the tricritical points that separates the second and first order regions in the $\mu-m_{K}$ plane, starts on the $m_{\pi}=0$ axis from the blue dot respectively at the $m_{K}^{TCP}=137.2$ MeV in (a) for the QMVT model and $m_{K}^{TCP}=242.6$ MeV in (b) for the RQM model. The blue dash line in the $m_{\pi}-m_{K}$ plane, depicts the strange chiral limit $h_{y}=0$ and the $h_{y}$ is negative in the shaded unphysical region. The black dash vertical line of crossover transition at the physical point, ends at the critical end point with $\mu_{CEP}=288.03\text{ MeV}$ in (a) for the QMVT model and $\mu_{CEP}=243.29\text{ MeV}$ in (b) for the RQM model. The vertical red line depicts the first order transition.
@@
The variation of the $m_{\sigma}$ with respect to the temperature is very smooth in the Fig. 3 of the QMVT model when $\beta=1.0\text{ and }0.5$ while the $m_{\sigma,c}$ RQM model plot in the Fig. 3 is sharper for the $\beta=0.5$ though being quite smooth at the physical point $\beta=1.0$. The respective temperature plots of the QMVT model $m_{\pi}$ and the RQM model $m_{\pi,c}$ in the Fig. 3 and 3, begin from their reduced value for $\beta<1$ and become degenerate respectively with the $m_{\sigma}$ and $m_{\sigma,c}$ temperature variations for the corresponding $\beta$ after the chiral transition takes place. The vacuum $\pi$ curvature mass from where the $m_{\pi,c}$ plots begin in the Fig. 3 are,  $m_{\pi,c}=135.98,68.75,50.17\text{ and }20.69$ in respective order for the $\beta=1.0,0.5,0.3642\text{ and }0.15$. The $m_{\sigma}$ becomes zero in its temperature variation in the Fig. 3 when the $\beta_{cep}=0.2028$ signaling identification of the critical end point in the QMVT model. The signature identification of the critical end point results in the RQM model when the curvature mass of the $\sigma$ meson $m_{\sigma,c}$ becomes zero in its temperature variation in the Fig. 3 for the $\beta_{cep}=0.3642$. For very small $\beta=0.15$ close to the chiral limit, the QMVT model $m_{\sigma}$ temperature plot in the Fig. 3 and the RQM model $m_{\sigma,c}$ temperature plot in Fig. 3, show a large mass gap that corresponds to the strong first order chiral transition.
@@ III.3 Comparing various aspects of Columbia plots in the QMVT and RQM model for different $\bf{m_{\sigma}}$
The QMVT model tricritical line for the $m_{\sigma}=530$ MeV that starts from the $m_{K}^{TCP}=124.7,\text{ at }\mu=0$ MeV in the Fig. 5, shows a good saturation trend similar to the $m_{\sigma}=400$ MeV case when chemical potential ($\mu\equiv 283$ MeV) is significantly higher. The average slope of the above line is only .066 between the points $(m_{K},\mu)=(260,266.82)$ MeV and $(m_{K},\mu)=(500,282.77)$ MeV and the slope becomes zero for the $m_{K}=450-500$ MeV range as $\mu=282.85$ MeV for the $m_{K}=450$ MeV, becomes only $\mu=282.77$ MeV for the $m_{K}=500$ MeV. When $m_{K}=550$ MeV, $\mu$ decreases slightly to $\mu=281.81$ MeV and the slope becomes slightly negative. In complete contrast to the QMVT model, the tricritical line of RQM model for $m_{\sigma}=530$ MeV starting from the $m_{K}^{TCP}=235.7,\text{ at }\mu=0$ MeV in Fig. 5, does not show a near saturation trend like the one that we get for the $m_{\sigma}=400$ MeV case. The slope 0.1852 between the points ($\mu=219.14,m_{K}=400.0$) MeV and ($\mu=228.4,m_{K}=450.0$) MeV on the tricritical line in the Fig. 5, decreases to the 0.1478 when one goes from the point ($\mu=228.4,m_{K}=450.0$) MeV to the point ($\mu=235.79,m_{K}=500.0$) MeV but it increases and becomes 0.1566 between the point ($\mu=235.79,m_{K}=500.0$) MeV and ($\mu=243.62,m_{K}=550.0$) MeV. The critical end point marked by the solid blue arrow, gets located at the $\mu_{CEP}=288.03$ MeV for the $m_{\sigma}=400$ MeV and the $\mu_{CEP}=322.16$ MeV for the $m_{\sigma}=530$ MeV case in the QMVT model while in comparison the position of the critical end point is at the $\mu_{CEP}=243.29$ MeV for the $m_{\sigma}=400$ MeV and the $\mu_{CEP}=275.96$ MeV in the RQM model.
[…]
@@
Table 3: This table presents the quantities $m_{K}^{TCP},\ m_{s}^{TCP},\ m_{\pi}^{t},\ m_{ud}^{t},\ m_{\pi}^{c},\ m_{ud}^{c}$ and the $\beta_{cep}:m_{\pi,cep}^{*},m_{K,cep}^{*}$ calculated for $m_{\sigma}=400\text{ and }500$ MeV in the QMVT and RQM model. RQM model results obtained using the large $N_{c}$ standard U(3) ChPT input for $m_{\sigma}=530$ MeV, is also presented. Some of these quantities as obtained in the e-MFA:QM model FRG study for $m_{\sigma}=530$ MeV and the QM model s-MFA study for $m_{\sigma}=800$ MeV, are also presented for comparison. The second order chiral transition and the TCP is absent on the $m_{\pi}=0$ axis in the QM model s-MFA study. The last two rows show the critical pion (light quark) mass $m_{\pi}^{c}$ ($m_{ud}^{c}$) for three degenerate flavors at $\mu=0$ from different LQCD studies whose results are labeled with the corresponding fermion implementation and the number of time slices $N_{t}$.
@@
QMVT $400$ 137.2 8.06 98.15 2.05 78.45 1.31 0.2215 30.57 109.88
RQM $400$ 242.6 25.53 168.39 6.13 134.16 3.84 0.3795 52.38 188.25
QMVT $530$ 124.7 6.64 90.03 1.72 72.12 1.1 0.2028 27.98 100.58
RQM $530$ 235.7 24.08 160.75 5.58 128.38 3.51 0.3642 50.26 180.64
RQM : ChPT-II $530$ 239.58 23.47 161.35 5.44 128.04 3.45 0.36778 50.9 182.96
e-MFA:QM-FRGResch $530$ 169 - 110 3.0 86 - 0.20 27.6 99.2
s-MFA:QMSchaefer:09 $800$ No TCP 177 - 150 - 0.488 67.34 242.05

### v2

@@ III.1 QMVT versus RQM model order parameters and its derivatives in the path to chiral limit
It is worthwhile to compare the $\beta_{cep}$ for the case of the $m_{\sigma}=400$ MeV also, though the order parameter and its derivative plots are not shown. The critical end point fraction in the QMVT model is found at the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{QMVT})}=0.2215\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(30.57,~109.88)$ MeV when the $m_{\sigma}=400$ MeV while the similar cep in the RQM model is found at the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{RQM})}=0.3795\text{ for }(m_{\pi,cep}^{*},m_{K,cep}^{*})=(52.38,~188.25)$ MeV. It is known that the strength of the chiral transition becomes stronger for smaller scalar $\sigma$ meson masses. The above feature gets verified as we are finding the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{QMVT})}=0.2215>\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{QMVT})}=0.2028$ in the QMVT model while we find the $\beta_{cep}|_{(m_{\sigma}=400\text{ MeV}:\text{RQM})}=0.3795>\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{RQM})}=0.3642$ for the RQM model. The $\beta_{cep}$ computed using the large $N_{c}$ standard U(3) ChPT input set $\text{M}_{\eta}$-II in the RQM model gives slightly different $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{RQM};\text{M}_{\eta}-\text{II})}=0.36887$.
It is important to compare the QMVT model result of the present work with the similar result of the Ref. Resch where the effect of quark one-loop vacuum fluctuation was computed in the e-MFA:QM Model FRG work under the LPA by switching off the quantum and thermal fluctuations of the mesons in the functional renormalization group flow equation. Recall that in the Ref. Resch , they reduce the non-strange and strange direction chiral symmetry breaking source strengths defined as $j_{x}\text{ and }j_{y}$ in their work by tuning a single parameter $\alpha=(0,0.04,0.17,1)$ but corresponding to each of the above reduction in the $\alpha$, they respectively adjust the initial action by changing the scale $\Lambda=(1143,1100,1000,700)\text{ MeV }$ such that the $f_{\pi}$ always remains fixed and the spontaneous chiral symmetry breaking does not get lost. The crossover transition for the $\alpha=1$ at the physical point, becomes second order for the critical $\alpha_{c}=.04$ and one gets the first order transition for the $\alpha_{c}<.04$ when the $m_{\sigma}=530$ MeV in their work. The parameter $\alpha$ of their study gives the related parameter $\beta=\sqrt{\alpha}$ in our work. Comparing our QMVT model results with those of the Ref. Resch , we see that the second order $Z_{2}$ critical point in their e-MFA:QM model study occurs at $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{eMFA;QM-FRG})}=\sqrt{\alpha_{c}}=0.2$ for $(m_{\pi,cep}^{*},m_{K,cep}^{*})=(27.6,~99.2)$ MeV which is quite close to our QMVT model result of the $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{QMVT})}=0.2028$. Thus the smoothing effect of fermionic vacuum fluctuation on the chiral transition is significantly large (and similar) in the QMVT model as well as the e-MFA:QM model FRG study of Ref. Resch where the extent of first order transition region is small. The above softening effect is quite moderate in the RQM model study which has already been reported very recently in the Ref. vkt25 where one finds a noticeably larger extent of the first order region and quite a large $\beta_{cep}|_{(m_{\sigma}=530\text{ MeV}:\text{RQM})}=0.3642$ also. Here it is relevant to point out that when the Dirac’s sea contribution gets neglected under the s-MFA in the QM model, the Fig. (9) of the Ref. Schaefer:09 , shows that the crossover transition ends at the second order transition for the significantly larger fraction $\beta_{cep}=.488$ giving noticeably larger $\pi\text{ and }K$ masses $(m_{\pi,cep}^{*},m_{K,cep}^{*})=(67.34,~242.05)$ MeV where they have taken $m_{\sigma}=800\text{ MeV}$ in their calculation so that the SCSB is not lost. The Table (3) shows the comparison of different results for the chiral limit studies done in the QMVT model, RQM model, e-MFA:FRG QM model and the s-MFA QM model. The table shows the LQCD results also for the ready reference.
@@ III.2 Comparing the $m_{\sigma}$ towards the chiral limit
Figure 4:  Columbia plot with the $U_{A}$(1) anomaly is presented in the left (right) panel (a) ((b)) for the QMVT (RQM) model when $m_{\sigma}=400$ MeV. The $\mu-m_{K}$ ($m_{\pi}-m_{K}$) plane presents the chiral transition for the $m_{\pi}=0$ ($\mu=0$). The second order $Z(2)$ critical solid black line demarcating the crossover from the first order region, intersects the SU(3) symmetric $m_{\pi}=m_{K}$ green dash line respectively at the critical pion mass of $m_{\pi}^{c}=78.45$ and $m_{\pi}^{c}=134.16$ MeV in (a) for the QMVT model and (b) for the RQM model. It terminates at the terminal pion mass of $m_{\pi}^{t}=98.15$ MeV in (a) for the QMVT model and $m_{\pi}^{t}=168.39$ MeV in (b) for the RQM model. Solid blue line of the tricritical points that separates the second and first order regions in the $\mu-m_{K}$ plane, starts on the $m_{K}$ ($m_{\pi}=0$) axis from the blue dot respectively at the $m_{K}^{TCP}=137.2$ MeV in (a) for the QMVT model and $m_{K}^{TCP}=242.6$ MeV in (b) for the RQM model. The blue dash line in the $m_{\pi}-m_{K}$ plane, depicts the strange chiral limit $h_{y}=0$ and the $h_{y}$ is negative in the shaded unphysical region. The black dash vertical line of crossover transition at the physical point, ends at the critical end point with $\mu_{CEP}=288.03\text{ MeV}$ in (a) for the QMVT model and $\mu_{CEP}=243.29\text{ MeV}$ in (b) for the RQM model. The vertical red line depicts the first order transition.
Note that the curvature and pole masses are same in the QMVT model whereas the curvature mass of the $\pi(\sigma)$ meson $m_{\pi,c}(m_{\sigma,c})=135.98(376.74)\text{ MeV}$ is smaller (significantly smaller) than its pole mass of $m_{\pi}(m_{\sigma})=138(530)\text{ MeV}$ for $\beta=1$ in the RQM model due to the on-shell renormalization of the parameters. The QMVT model temperature variation of the $\sigma$ mass in the Fig. 3, begins from the same value $m_{\sigma}=530$ MeV for all the $\beta$ fractions whereas the RQM model curvature mass of the $\sigma$ meson in the Fig. 3 starts from the $m_{\sigma,c}=376.74,420.29,443.01\text{ and }474.75$ MeV respectively for the $\beta=1.0,0.5,0.3642\text{ and }0.15$. Thus we find that the vacuum curvature mass of the $\sigma$ meson in the RQM model, is lowest at the physical point and it increases successively when one moves towards the chiral limit after reducing the ($m_{\pi},m_{K}$) by a fraction $\beta$ of their physical point value. The behavior of the $m_{\sigma}$ in our chiral limit study for both the QMVT and RQM models, stands in contrast to its behavior in the e-MFA-FRG, QM model study of the Ref.Resch where in the aftermath of the successive increase in the scale $\Lambda=700,1000,1100\text{ and }1143$ corresponding to the successive decrease of the source strengths $j_{x}\text{ and }j_{y}$ governed by the $\alpha=\sqrt{\beta}=1.0,0.17,0.04\text{ and }0$, the $\sigma$ mass registers a significant successive decrease towards the chiral limit as evident from the Fig.2(c) of the Ref.Resch . Here, it is relevant to point out that the $\sigma$ mass in the e-MFA-FRG study is calculated from the curvature of the effective potential that gets evaluated by the FRG flow equation. Recall that the curvature masses of mesons are used for fixing the parameters in the QMVT model and this context of incorporating quark one-loop vacuum correction is analogous to that of the e-MFA-FRG study in the Ref.Resch where the vacuum value of the $m_{\sigma}$ decreases significantly towards the chiral limit while the QMVT model $m_{\sigma}$ does not change towards the chiral limit. In complete contrast of the above two scenarios, being lowest at the physical point and different from its pole mass $m_{\sigma}=530$ MeV, the vacuum curvature mass $m_{\sigma,c}$ in the on-shell renormalized RQM model increases towards the chiral limit.
@@
The variation of the $m_{\sigma}$ with respect to the temperature is very smooth in the Fig. 3 of the QMVT model when $\beta=1.0\text{ and }0.5$ while the $m_{\sigma,c}$ RQM model plot in the Fig. 3 is sharper for the $\beta=0.5$, though being quite smooth at the physical point $\beta=1.0$. The respective temperature plots of the QMVT model $m_{\pi}$ and the RQM model $m_{\pi,c}$ in the Fig. 3 and 3, begin from their reduced value for $\beta<1$ and become degenerate respectively with the $m_{\sigma}$ and $m_{\sigma,c}$ temperature variations for the corresponding $\beta$ after the chiral transition takes place. The vacuum $\pi$ curvature mass from where the $m_{\pi,c}$ plots begin in the Fig. 3 are,  $m_{\pi,c}=135.98,68.75,50.17\text{ and }20.69$ in respective order for the $\beta=1.0,0.5,0.3642\text{ and }0.15$. The $m_{\sigma}$ becomes zero in its temperature variation in the Fig. 3 when the $\beta_{cep}=0.2028$ signaling identification of the critical end point in the QMVT model. The signature identification of the critical end point results in the RQM model when the curvature mass of the $\sigma$ meson $m_{\sigma,c}$ becomes zero in its temperature variation in the Fig. 3 for the $\beta_{cep}=0.3642$. For very small $\beta=0.15$ close to the chiral limit, the QMVT model $m_{\sigma}$ temperature plot in the Fig. 3 and the RQM model $m_{\sigma,c}$ temperature plot in Fig. 3, show a large mass gap that corresponds to the strong first order chiral transition.
@@ III.3 Comparing various aspects of Columbia plots in the QMVT and RQM model for different $\bf{m_{\sigma}}$
Columbia plot studies are being actively pursued in different approaches in the light of the most recent lattice results. Here it is relevant to point out that apart from the cubic ‘t Hooft coupling term, the sixth order anomalous coupling (which becomes relevant pisarski24 ) term might also govern the change in the size of first order regions as shown in the Fig.1 of a very recent work on the Columbia plot in Ref.Giacosa . They have discussed a scenario where the extent of first order region is maximum when the sixth order anomalous coupling is zero and the first order region decreases corresponding to the increasing sixth order coupling strength while the cubic coupling strength decreases. They have shown a possibility where phase transition in the whole $m_{\pi}-m_{K}$ plane of the Columbia plot becomes second order, even though the $U_{A}(1)$ anomaly is still broken because of the large strength of the sixth order anomalous coupling term while the cubic coupling strength is zero. Putting the present work in comparative perspective, the size of the first-order regions also differ significantly between the two models, QMVT and RQM in the Fig.4 and Fig.5. Note that our work considers only the standard cubic coupling term $c$ which being same as in the QM model, does not change in the curvature mass based parameter fixing of the QMVT model and significantly reduced first order regions of the Fig.4 and Fig.5 are consequence of the very large smoothing effect of the quark one-loop vacuum fluctuation on the strength of the chiral transition. In contrast to the above, the large first order regions in the Fig.4 and Fig.5 of the RQM model Columbia plots are caused by the fact that the $U_{A}(1)$ anomaly strength $c$ which contains a condensate dependent part,  gets significantly enhanced when the meson self energies due to quark loops are evaluated using the meson pole masses and parameters are fixed on-sell. The smoothing effect of the quark one-loop vacuum fluctuation is moderate in RQM Model.
@@
Figure 7: Left panel (a) (Right panel (b)) shows the comparison of lines and regions in the vertical $\mu-m_{K}$, $m_{\pi}=0$ (horizontal $m_{\pi}-m_{K}$, $\mu=0$) planes of the three dimensional QMVT and RQM model Columbia plots for the $m_{\sigma}=400,\ 530$ MeV in the Fig.(4), Fig.(5) and Fig. (4), Fig.(5). QMVT and RQM model tricritical lines for the $m_{\sigma}=400\text{ and }530$ MeV, as marked and explained in the Fig.(a), start from the $m_{K}^{TCP}$ on the $m_{K}$ axis and demarcate the first order region above them on their left hand side from the second order region below them on their right hand side. QMVT and RQM model Z(2) critical lines for the $m_{\sigma}=400\text{ and }530$ MeV, as marked and explained in the Fig.(b), separating the crossover transition region above them from the first order region that lies below, intersect the SU(3) symmetric $m_{\pi}=m_{K}$ chiral limit line at the critical pion mass $m_{\pi}^{c}$ and terminate on the strange chiral limit line satisfying the $h_{y}=0=h_{y0}$, at the terminal pion mass $m_{\pi}^{t}$.
The QMVT model tricritical line for the $m_{\sigma}=530$ MeV that starts from the $m_{K}^{TCP}=124.7,\text{ at }\mu=0$ MeV in the Fig. 5, shows a good saturation trend similar to the $m_{\sigma}=400$ MeV case when chemical potential ($\mu\equiv 283$ MeV) is significantly higher. The average slope of the above line is only .066 between the points $(m_{K},\mu)=(260,266.82)$ MeV and $(m_{K},\mu)=(500,282.77)$ MeV and the slope becomes zero for the $m_{K}=450-500$ MeV range as $\mu=282.85$ MeV for the $m_{K}=450$ MeV, becomes only $\mu=282.77$ MeV for the $m_{K}=500$ MeV. When $m_{K}=550$ MeV, $\mu$ decreases slightly to $\mu=281.81$ MeV and the slope becomes slightly negative. In complete contrast to the QMVT model, the tricritical line of RQM model for $m_{\sigma}=530$ MeV starting from the $m_{K}^{TCP}=235.7,\text{ at }\mu=0$ MeV in Fig. 5, does not show a near saturation trend like the one that we get for the $m_{\sigma}=400$ MeV case. The slope 0.1852 between the points ($\mu=219.14,m_{K}=400.0$) MeV and ($\mu=228.4,m_{K}=450.0$) MeV on the tricritical line in the Fig. 5, decreases to the 0.1478 when one goes from the point ($\mu=228.4,m_{K}=450.0$) MeV to the point ($\mu=235.79,m_{K}=500.0$) MeV but it increases and becomes 0.1566 between the point ($\mu=235.79,m_{K}=500.0$) MeV and ($\mu=243.62,m_{K}=550.0$) MeV. The critical end point marked by the solid blue arrow, gets located at the $\mu_{CEP}=288.03$ MeV for the $m_{\sigma}=400$ MeV and the $\mu_{CEP}=322.16$ MeV for the $m_{\sigma}=530$ MeV case in the QMVT model while in comparison the position of the critical end point is at the $\mu_{CEP}=243.29$ MeV for the $m_{\sigma}=400$ MeV and the $\mu_{CEP}=275.96$ MeV for the $m_{\sigma}=530$ MeV in the RQM model.
@@
Table 3: This table presents the quantities $m_{K}^{TCP},\ m_{s}^{TCP},\ m_{\pi}^{t},\ m_{ud}^{t},\ m_{\pi}^{c},\ m_{ud}^{c}$ and the $\beta_{cep}:m_{\pi,cep}^{*},m_{K,cep}^{*}$ calculated for $m_{\sigma}=400\text{ and }500$ MeV in the QMVT and RQM model. RQM model results for $m_{\sigma}=530$ MeV obtained using the inputs of large $N_{c}$ standard U(3) ChPT termed as input set $\text{M}_{\eta}$-II in Ref.vkt25 , is also presented. Some of these quantities as obtained in the e-MFA:QM model FRG study for $m_{\sigma}=530$ MeV and the QM model s-MFA study for $m_{\sigma}=800$ MeV, are also presented for comparison. The second order chiral transition and the TCP is absent on the $m_{\pi}=0$ axis in the QM model s-MFA study. The last two rows show the critical pion (light quark) mass $m_{\pi}^{c}$ ($m_{ud}^{c}$) for three degenerate flavors at $\mu=0$ from different LQCD studies whose results are labeled with the corresponding fermion implementation and the number of time slices $N_{t}$.
@@
QMVT 400400 137.2 8.06 98.15 2.05 78.45 1.31 0.2215 30.57 109.88
RQM 400400 242.6 25.53 168.39 6.13 134.16 3.84 0.3795 52.38 188.25
QMVT 530530 124.7 6.64 90.03 1.72 72.12 1.1 0.2028 27.98 100.58
RQM 530530 235.7 24.08 160.75 5.58 128.38 3.51 0.3642 50.26 180.64
RQM : $\text{M}_{\eta}$-II vkt25 530530 239.58 23.47 161.35 5.44 128.04 3.45 0.36887 50.9 182.96
e-MFA:QM-FRGResch 530530 169 - 110 3.0 86 - 0.20 27.6 99.2
s-MFA:QMSchaefer:09 800800 No TCP 177 - 150 - 0.488 67.34 242.05